OpenStudy (anonymous):

The grades on the last math exam had a mean of 88%. Assume the population of grades on math exams is known to be normally distributed with a standard deviation of 6. What percent of students earn a score between 72% and 90%?

OpenStudy (anonymous):

\[P(72<X<90)=P\left(\frac{72-88}{6}<\frac{X-88}{6}<\frac{90-88}{6}\right)=P(-2.67<Z<0.33)\]